### v1

# Néel-Vector-Orientation Induced Intrinsic Half-Metallicity in Two-Dimensional Altermagnets

Abstract
Altermagnets combine zero net magnetization with giant spin splitting, enabling spin‑polarized transport without strong spin–orbit coupling (SOC). Deterministically selecting the conducting spin channel, however, requires breaking the $90^{\circ}$ rotation–time‑reversal antisymmetry ($C_{4z}\mathcal{T}$). Using standard axial‑vector transformation rules as preliminaries, we show that in monolayer Ta2TeSeO this can be achieved naturally and tuned in a symmetry‑efficient way by rotating the Néel vector. Without considering Néel vector, Ta2TeSeO have one pair of mirror-protected spin-polarized Weyl point in each spin channel. Aligning the Néel vector along the crystallographic $x$ or $y$ direction breaks the mirror symmetry $M_{x}$ or $M_{y}$, induce selective mirror symmetry breaking and preserving keeps one spin sector gapless, and opens a gap in the opposite spin, yielding fully spin‑polarized transport. The C2z symmetry breaking make the preserved two Wely points not equivalent, change the half-semimetal into half-metallic states. The same orientation‑selective symmetry reduction applies to lattice vibrations, implying phonon chirality splitting. Owing to the near‑degenerate in‑plane anisotropy, reversible zero‑moment switching is achievable with minute in‑plane strain or weak magnetic fields, and the lattice coupling suggests control by circularly polarized light. The mechanism extends to other two-dimensional decorated Lieb altermagnets with horizontal mirror $M_{z}$ absent, providing a general, low‑power route to spin filtering and logic.
Magnetic space group (MSG) theory extends crystallographic symmetry by treating time reversal on equal footing with spatial operations and thereby provides the natural language to describe symmetry reduction caused by magnetic ordering. In collinear antiferromagnets that lack $P\mathcal{T}$ or any other antiunitary constraint enforcing Kramers degeneracy, the MSG admits the nonrelativistic spin–crystal‑rotation symmetries characteristic of altermagnets, so that the two spin subspaces can become inequivalent at a given $\mathbf{k}$ and bands are spin split even when spin–orbit coupling (SOC) is negligible [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Once the magnetic axis (Néel vector) is specified as an axial vector, the way it transforms under mirrors and composite rotations further selects which operations survive as unitary versus magnetic ones, lowering the magnetic point group (e.g., $C_{4z}\mathcal{T}\to C_{2}\mathcal{T}$) and enabling additional symmetry‑lowering effects such as chirality splitting and pronounced $C_{2}$ anisotropies in the electronic and phonon band structure [11, 12, 13, 14, 15].
Realizing a single‑spin Fermi surface in a compensated magnet is nontrivial[16, 17, 18, 19, 20, 21, 22, 23, 24], and the most symmetry‑efficient parent phases are Weyl/Dirac semimetals rather than ordinary metals. In a semimetal, symmetry‑protected point (or line) nodes at EF can be selectively gapped by targeted symmetry lowering. In contrast, extended Fermi surfaces of trivial metals typically survive small symmetry changes and cannot be cleanly removed without large band reconstructions. Semimetals thus provide switchability and minimal energetic cost. Metals also have its unreplaceable advantagement over semimetal, they provide a finite density of states at EF for robust current drive, low contact resistance, and efficient spin injection. Combining the two, using symmetry to convert a half‑semimetal into a half‑metal is therefore ideal: breaking the protecting mirror/rotation both permits a mass term at one member of a symmetry‑related Weyl pair [25] and shifts the two nodes in opposite energy-level directions [26], so that one Weyl cone is pushed above EF while the other one is pulled below. The gaped spin channel becomes insulating, the preserved spin channel retains a finite single‑spin Fermi surface, and half-metallic state is achieved. This symmetry-guided pathway underlies the previously reported half-semimetallic transport in 2D altermagnets such as CrO and MoO [19, 27, 28, 29], where unequal strain displaces and annihilates spin-polarized Weyl points, thereby driving transitions between half-metallic and half-semiconducting states. Can intrinsic half-metallic states exist in antiferromagnets without any external strain? Here we demonstrate that such a conversion can be achieved intrinsically through magnetic-axis–controlled symmetry breaking, enabling deterministic spin-channel switching without invoking lattice distortions.
Figure 1: (a) Mirror-symmetry operation $M_{x}$ with respect to a plane normal to the $x$ axis applied to magnetic moments for different Néel-vector orientations $\mathbf{n}(x,y,z$, and $d\equiv(110)$ ). Black arrows indicate the moment directions, while red circular arrows depict the direction of the equivalent loop current used to visualize the moment, blue lines/rectangles mark the mirror plane. Owing to the axial-vector character of the magnetic moment, $M_{x}$ leaves the $x$ component unchanged but reverses the components parallel to the mirror plane, i.e., $\left(m_{x},m_{y},m_{z}\right)\rightarrow\left(m_{x},-m_{y},-m_{z}\right)$. (b) Rotation-symmetry operation $C_{2z}$ applied to magnetic moments for different Néel-vector orientations $\mathbf{n}(x,y,z$, and $d\equiv(110)$ ).
In this work, we revisit the standard symmetry operations on axial vectors (magnetic moments) within an orientation-resolved framework in which the Néel vector acts as an intrinsic generator of MSG reduction in altermagnets. Because an axial vector reverses different components under the mirror or rotation operations, fixing the magnetic axis converts selected unitary mirrors $m$ into magnetic mirrors $m\mathcal{T}$ or eliminates them entirely. In Lieb and decorated Lieb lattice [30, 19, 31, 28, 32, 33, 34, 28, 35] with horizontal mirror $M_{z}$ absent and spin-polarized Weyl points, the machinism decides which high‑symmetry lines retain mirror‑protected single‑spin crossings and which acquire SOC‑allowed mass terms[36]. Monolayer Ta2TeSeO provides a particularly clean two-dimensional platform for this mechanism. Theory predicts giant momentum-dependent spin splitting and spin-polarized, anisotropic Weyl nodes pinned near the Fermi level, supplying the band-geometry needed to reconcile integer filling with gapless transport. Equally important for control, the in-plane easy-axis anisotropy along $x$ or $y$ is degenerate, so rotating the Néel vector is energetically inexpensive. We show that aligning the magnetic axis along the crystallographic $x$ or $y$ direction breaks the mirror symmetry $M_{y}$ or $M_{y}$, leaving one spin sector gapless with half-metal-like dispersion while opening a gap in the other. The magnetic axis along $x$ or $y$ also reduce the $C_{2z}$ symmetry into $C_{2z}\mathcal{T}$ and thus break the equivalence of the left two Weyl points, thus realizing intrinsic half-metallic states in the altermagnet.
Because the easy-plane anisotropy is tiny, minute in-plane strain or weak magnetic fields suffice to reversibly switch the conducting spin channel. Coupling between the lattice and the Néel vector further suggests control by circularly polarized light [37, 38, 39]. The same orientation-controlled symmetry breaking generalizes beyond Ta2TeSeO and applies even to altermagnets lacking Weyl cones, offering a low-power route to spin filtering and logic that complements strain-driven and SOC-based control paradigms.
To clarify how the Néel-vector $\mathbf{n}$ orientation affects the surviving magnetic symmetries, we briefly recall the transformation rules of an axial vector under mirror and twofold rotation operations (Fig. 1). In the absence of SOC and without fixing $\mathbf{n}$, a mirror relates two atoms with the same magnetic moments and hence enforces their equivalence. Once $\mathbf{n}$ is specified, however, a mirror acts differently on magnetic moment versus polar vectors such as atomic displacement and may no longer relate the two states. The resulting symmetry reduction enables chirality splitting and other $C_{2}$-type anisotropies in the electronic and phonon spectra.
As an example, let the mirror plane be yzyz ($M_{x}$). For an axial vector $\mathbf{S}=(S_{x},S_{y},S_{z})$, the transformation is
$M_{x}:\ (S_{x},S_{y},S_{z})\ \mapsto\ (S_{x},-S_{y},-S_{z}),$ (1)
where the extra minus sign for the in-plane components originates from $\det M_{x}=-1$. This is conveniently visualized by the loop-current picture [15] of a magnetic moment in Fig. 1(a): a loop whose normal is along $x$ is unchanged by $M_{x}$, while loops with normals along $y$ or $z$ reverse their circulation. Consequently:
  • for $\mathbf{n}\parallel x$, the unitary mirror $M_{x}$ survives;
  • for $\mathbf{n}\parallel y$ or $z$, $M_{x}$ inverts the moment and must be accompanied by time reversal to be a symmetry (magnetic mirror $M_{x}\mathcal{T}$);
  • for $\mathbf{n}\parallel d\equiv(110)$, $M_{x}$ maps $d\to[1\bar{1}0]$, i.e., rotates the in-plane axis by $90^{\circ}$ within the layer.
Rotations act on axial and polar vectors identically, so
$C_{2z}:\ (S_{x},S_{y},S_{z})\ \mapsto\ (-S_{x},-S_{y},S_{z}).$ (2)
For an in-plane Néel vector, $C_{2z}$ reverses the moment and must be combined with time reversal to remain a symmetry, thus the composite antiunitary $C_{2z}\mathcal{T}$ is preserved when $\mathbf{n}\perp z$, whereas it is generally broken for $\mathbf{n}\parallel z$ (because $\mathcal{T}$ flips $S_{z}$).
In a Janus decorated Lieb lattice with spin-polarized Weyl points, $M_{z}$ is naturally broken, $M_{x}$ and $M_{x}$ can be selectively broken by an in-plane Néel vector. Combining this with the reduction from the fourfold composite $C_{4z}\mathcal{T}$ to $C_{2z}\mathcal{T}$ when $\mathbf{n}$ is fixed along $x$ or $y$, two direct consequences can be implied:
  1. 1.
    Weyl-pair in different spin-channel inequivalence. Nodes related by $90^{\circ}$ rotation are no longer symmetry-locked once only $C_{2z}\mathcal{T}$ remains, so they may acquire different energies or gaps.
  2. 2.
    Selective gapping. Symmetry now permits a mass term in one spin sector while forbidding it in the other, leaving one sector gapless with half-metal-like dispersion and gapping the opposite sector.
  3. 3.
    Weyl-pair in same spin-channel inequivalence. Nodes related by $180^{\circ}$ rotation are no longer symmetry-locked since the mirror symmetry breaking and $C_{2z}$ symmetry breaking, so they may acquire different energies or gaps. This can change semimetal to metal.
Taken together, the mirror-selection and the $C_{2z}\mathcal{T}$ reduction provide a compact, orientation-resolved mechanism for symmetry-controlled spin-channel selection and chirality splitting, as we will demonstrate explicitly for monolayer Ta2TeSeO.
Figure 2: (a) Top and side views of monolayer Ta2TeSeO. Yellow, red, blue, orenge, and green spheres denote O atoms, spin-up-polarized Ta atoms, and spin-down-polarized Ta atoms, Se atoms, and Te atoms, respectively. (b) The first BZ of Ta2TeSeO, with all Weyl points are shown. The green and blue areas are the two irreducible parts of the first BZ when SOC effect is not taken into account. (c) The first BZ of Ta2TeSeO. The four areas in different color are the four irreducible parts of the first BZ when SOC effect is taken into account. (d) The band structure without SOC are shown. The up- and down-spin channels are depicted in red and blue, respectively.
We performed first-principles calculations within density functional theory as implemented in the Vienna ab initio Simulation Package (VASP) [40, 41]. The exchange–correlation potential was treated using the generalized-gradient approximation supplemented by a Hubbard correction (PBE+U) [42] for the Ta-$d$ orbitals, with an effective $U$ value of 3.2 eV [43]. To avoid spurious interlayer interactions in the out-of-plane direction, a vacuum spacing larger than 20 Å was introduced. The plane-wave basis was expanded up to a kinetic energy cutoff of 600 eV. Brillouin-zone sampling was carried out with a Monkhorst–Pack grid denser than $2\pi\times 0.033$ Å-1. The electronic self-consistency criterion was set to $10^{-6}$ eV, and atomic positions were relaxed until the residual forces on each atom were below 0.001 eV/Å. To assess the dynamical stability of the structures, phonon spectra were calculated using a $3\times 3\times 1$ k-points mesh with the PHONOPY code [44].
The Heisenberg exchange parameters were calculated by using the magnetic force theorem [45], which is implemented in the full-potential linear muffin-tin orbital (FP-LMTO) code RSPt[46]. The following effective spin Hamiltonian is then adopted for classical Monte Carlo simulations
$\hat{H}=-\sum_{i\neq j}J_{ij}\vec{S}_{i}\cdot\vec{S}_{j}-\sum_{i}K_{i}\left(S_{i}^{z}\right)^{2},$ (3)
where $J_{ij}$ is the Heisenberg exchange interaction between site $i$ and $j$, $\vec{S}$ is three-dimensional vector representing the magnetic moment at a given site, and $K$ denotes the strength of single ion magnetocrystalline anisotropy energy (MAE). The critical temperature of the magnetic phase transition was obtained by performing classical Monte-Carlo simulations using the UppASD package[47]. Periodic boundary conditions were applied in two dimensions ($xy$ plane), with a $80\!\times\!80\!\times\!1$ supercell (containing 12800 Ta atoms), an annealing process was simulated, starting at 500 K and gradually cooling to 0 K with 25000 Monte-Carlo steps performed at each temperature. For comparison, we used the mean-field approximation (MFA)[48]: $T_{N}=\frac{2J_{0}}{3k_{B}}$, where $J_{0}=\sum_{j}J_{0j}$ is the sum of the calculated Heisenberg exchange constants.
Figure 3: (a) Calculated exchange parameter $J_{ij}$ (black) and specific heat (blue) in Ta2TeSeO. Positive and negative exchange parameters correspond to FM and AFM couplings. The first-, second- and third-nearest-neighbor Ta-Ta pairs are illustrated by red, green, and cyan arrows in the inset plot. Dashed horizontal and vertical lines represent the zero exchange coupling strength and the phase transition temperature, respectively. (b) The phonon spectra along the high symmetry path in the first Brillouin zone.
Our calculated structure and magnetic configuration of Ta2TeSeO are shown in Fig. 2(a). The unit cell contains one Se, one Te, one O atom and two antiferromagnetically coupled Ta atoms ($\sim 0.73\mu B$). Calculated inter-site magnetic exchange parameters are shown in Fig. 3 (a), with the first nearest-neighbor couplings to be AFM (-7.31 meV), and the second and third are FM (1.13 meV and 2.38 meV), respectively. These results are fully consistent with the AFM ground state, as illustrated by the red, green, and cyan arrow on the Ta lattice shown in the inset. Using the calculated values $J_{ij}$, we determined the Néel temperature ($T_{N}$) of AFM Ta2TeSeO via both MFA and classical Monte Carlo simulations, yielding 346.9K346.9\ $\text${K} and 235K235\ $\text${K}, respectively. The latter is illustrated by the specific heat $C_{V}$ as a function of temperature in Fig. 3 (a), which serves a critical indicator due to the zero net magnetic moment in the AFM system. The in‑plane lattice parameter is about 4.22 Å. All Ta atoms lie in one plane, sandwiched by Se and Te atomic layers, showing a Janus structure. An oxygen atom is situated at the middle atomic plane close to Ta layer, and slightly buckled. In the nonmagnetic limit the monolayer adopts the symmorphic space group $P4mm$ (No. 99), whose point group is $C_{4v}$ generated by $C_{4z}$ and seveal vertical mirror symmetry. The horizontal mirror Mz is absent due to the Janus polarity. Dynamical stability are confirmed by phonon calculations, shown in Fig. 3 (b).
In the SOC-free limit, Ta2TeSeO factorizes into two spin blocks, and each block realizes a pair of symmetry-pinned Weyl cones on the Brillouin-zone boundary. The minimal irreducible high symmetry path and corrosponding band structure are shown in Figs. 2 (b) and (d), respectively. The pinning is enforced by unitary in-plane mirrors: along the $M_{3}-M_{1}$ and $M_{1}-M_{2}$ lines, the little group contains $M_{x}$ and $M_{y}$, respectively. Within a single-spin block the two crossing Bloch states transform according to different one-dimensional irreducible representations of the mirror, so the only $2\times 2$ mass term $m\sigma_{y}$ that could hybridize them [49],
$H_{0}(\mathbf{q})=v_{\|}q_{\|}\sigma_{x}+v_{\perp}q_{\perp}\sigma_{z},\quad M:\sigma_{y}\mapsto-\sigma_{y},$ (4)
is odd under the corresponding mirror and therefore forbidden by symmetry[50]. This $k\cdot p$ form makes explicit that a mirror‑invariant little group excludes the term $m\sigma_{y}$.
Turning on SOC, $\lambda\mathbf{L}\cdot\mathbf{S}$, couples spin to the crystalline orbital basis, and the AFM order selects a Néel vector $\mathbf{n}$ (an axial vector). Whether the mirror that protected the single-spin Weyl cone survives now depends on $\mathbf{n}$. Under a reflection $M_{x}$ (plane $yz$ ), an axial vector transforms as $\left(n_{x},n_{y},n_{z}\right)\mapsto\left(n_{x},-n_{y},-n_{z}\right)$. Thus, $\mathbf{n}\|\hat{x}$ preserves the unitary $M_{x}$ but breaks $M_{y}$. Conversely, $\mathbf{n}\|\hat{z}$ breaks both $M_{x}$ and $M_{y}$ as unitary symmetries (at most their antiunitary magnetic counterparts $M_{x,y}\mathcal{T}$ remain). The crucial point is that antiunitary operations do not provide a good quantum number that distinguishes the two crossing branches on a mirror-invariant line. Hence, once the corresponding unitary mirror is lost, the hybridization $m\sigma_{y}$ becomes symmetry-allowed. In an SOC language, the amplitude obeys
[…]
opening a gap $\Delta=2|m|$. Thus, by selecting the Néel‑axis orientation, one mirror remains unitary and protects the crossing in one spin sector, while the other mirror becomes magnetic (or absent) and permits a gap in the opposite sector, exactly the selective gapping mechanism resolved in Fig. 4.
Figure 4: Spin-projected electronic band structures of Ta2TeSeO for magnetic axis aligned along (a) the out-of-plane z direction, (b) the in-plane x direction, and (c) the in-plane diagonal (110) direction.
Selective mirror breaking explains why one set of nodes is gapped. What remains to explain is why the other pair of nodes, protected by the surviving mirror, are no longer pinned to the same energy, as shown in Fig. 4(b). The two surviving nodes (on $X\!-\!M_{1}$ and $X\!-\!M_{3}$) cease to be isoenergetic because fixing an in–plane x-direction Néel axis breaks the unitary twofold rotation $C_{2z}$. In the parent $C_{4v}$ phase, $C_{2z}$ maps $(\pi,k_{y})\!\mapsto\!(\pi,-k_{y})$ (modulo a reciprocal vector), enforcing $E(\pi,k_{y})=E(\pi,-k_{y})$ and hence equivalence of the two cones. With $\mathbf{n}\!\perp\!z$, only the antiunitary $C_{2z}\mathcal{T}$ may remain, it flips spin and imposes no constraint within a fixed spin block, so the nodes along $X\!-\!M_{1}$ and $X\!-\!M_{3}$ can shift to different energies. For the protected spin sector a minimal $k\!\cdot\!p$ reads
$H_{i}(\mathbf{q})$ $=v_{i,\parallel}q_{\parallel}\sigma_{x}+v_{i,\perp}q_{\perp}\sigma_{z}+m_{i}\sigma_{y}+\delta_{i}\sigma_{0},$ (6)
$\qquad i=1\ (X\!-\!M_{1}),3\ (X\!-\!M_{3}).$
mirror selection fixes $m_{1}=m_{3}=0$ for the protected sector, whereas breaking unitary $C_{2z}$ lifts the constraint $\delta_{1}=\delta_{3}$ and allows independent scalar shifts $\delta_{1,3}$, i.e., $E_{W}^{(1)}-E_{W}^{(3)}=\delta_{1}-\delta_{3}\neq 0$, so the two cones are no longer rotationally equivalent. This inequivalence enables the half–semimetal $\rightarrow$ half–metal conversion: by small symmetry‑compatible tuning (strain, gating) one cone is pushed above $E_{F}$ while the other is pulled below.
Figure 5: Symmetry–selected spin-polarized trasport under different Néel–vector orientations. Schematic distribution of spin–resolved Weyl cones on the Brillouin–zone boundary when the Néel vector is aligned (a) along $x$, (b) along $y$, (c) along $z$, and (d) along the in–plane diagonal $d\!\equiv\!(110)$. Red and orange colors denote the occupied and unocuppied states in spin-up channel, blue and cyan colors denote the occupied and unocuppied in spin-down channel.
The four representative configurations in Fig. 5 illustrate how the Néel-vector orientation governs symmetry and, consequently, the nature of spin-polarized transport. When $\mathbf{n}$ lies along $x$ or $y$, one in-plane mirror remains unitary while the orthogonal one becomes magnetic, preserving gapless Weyl cones in a single spin channel and producing half-metallic transport with opposite spin polarization for $\mathbf{n}\!\parallel\!x$ and $\mathbf{n}\!\parallel\!y$. For $\mathbf{n}\!\parallel\!z$, both in-plane mirrors are broken but $C_{2z}$ is retained, so both spin sectors open equal gaps, forming a compensated insulating state. For $\mathbf{n}$ along the diagonal $d\!\equiv\!(110)$, in-plane mirrors are broken and unitary $C_{2z}$ is also lost, leading to energy-inequivalent gaped states within each spin sector.
In summary, we introduce a Néel-vector–oriented framework for MSG reduction that intrinsically realizes half-metallicity in two-dimensional altermagnets. Using monolayer Ta2TeSeO as a prototype, we show that without SOC, each spin subspace hosts a pair of mirror-protected Weyl points. Fixing the Néel vector along $x$ ($y$) preserves the unitary mirror $M_{x}$ ($M_{y}$) while converting its orthogonal partner into a magnetic mirror, allowing a SOC-induced mass term only in one spin sector while keeping the other gapless. Moreover, breaking the unitary $C_{2z}$ symmetry (leaving only $C_{2z}\mathcal{T}$) lifts the equivalence between the two surviving Weyl cones, driving an intrinsic half-semimetal to half-metal transition and producing a single-spin Fermi surface. This orientation-selective mechanism also applies to lattice vibrations, implying phonon chirality splitting. Because the in-plane anisotropy is nearly degenerate, minute strain or weak magnetic fields can reversibly switch the conducting spin channel without net magnetization, while coupling to the lattice further suggests control via circularly polarized light. The same principle generalizes to other two-dimensional decorated Lieb altermagnets with horizontal mirror $M_{z}$ absent (e.g., V2SeSO, Nb2SeTeO), offering a universal, symmetry-efficient route to low-power spin filtering and logic.
Acknowledgements.
We would like to thank Shandong Institute of Advanced Technology and National Supercomputing Center (Shuguang) for providing computational resources. This work is supported by the National Natural Science Foundation of Shandong Province(Grant No. ZR2024QA040). Xin Chen thanks China scholarship council for financial support (No. 201606220031). Duo Wang acknowledges financial support from the Science and Technology Development Fund from Macau SAR (Grant No. 0062/2023/ITP2 and No. 0016/2025/RIA1) and the Macao Polytechnic University (Grant No. RP/FCA-03/2023). Biplab Sanyal acknowledges financial support from Swedish Research Council (grant no. 2022-04309), STINT Mobility Grant for Internationalization (grant no. MG2022-9386) and DST-SPARC, India (Ref. No. SPARC/2019-2020/P1879/SL).